# LLM-Enhanced User-Item Interactions: Leveraging Edge Information for Optimized Recommendations

### v1

@@ 4.2.2. Ablation Studies
  • LLM-NoGHIP does not include graph or historical information to embed prompts for pre-training, but only uses simple users’ review information to evaluate the impact of complex prompts on model performance.
We pre-trained and fine-tuned each baseline model separately, and then compared it with our complete model. These pre-training and fine-tuning experimental settings are consistent and conducted on the same dataset to ensure comparability of results. Table  2 shows the specific contribution of each component to the overall performance of the model. For example, the performance of LLM-NoPretrain is significantly lower than that of the complete model. This implies that using recommendation related graph data and natural language data for pre-training plays a crucial role in improving model performance. Similarly, the results of LLM-NoFineTune demonstrate the importance of fine-tuning. Subsequently, by comparing the performance of LLM-NoGKIA, and LLM-NoGHIP with that of the complete model, we find that the addition of graph connection information in attention calculation and complex prompts containing second-order relationships is crucial for improving the performance of recommendation systems.
4.2.3. Study on Different Prompt Structures
This experiment aims to answer: What is the contribution of text data that integrates second-order relationships and item background information to recommendation models? To answer this question, we chose the AM-Toys dataset and designed the following experimental models:
@@ 4.2.3. Study on Different Prompt Structures
  • Without 2-Order or Item: The model does not contain any second-order relationship information or item background information.
@@
Figure  4 shows that the model that uses prompt sentences of complete information (with the second-order relationship) performs best over all the performance indicators. The performances of the ”Without second-order relationship” model are lower than that of the complete model. As can be seen, second-order relationship information is an essential component of graph connectivity. Similarly, the ”Without Item” model performs poorly, highlighting the importance of natural language background information in enhancing recommendation systems.
@@ 4.2.4. Study on Different Attention Injection Ways
Figure 5 shows that the model using our graph attentive LLM method exhibits the best performance. Our method not only considers the direct connections between nodes but also the spatial relationships (i.e., the shortest connected path) between nodes in the graph. We compared our method with regular attention mechanisms, and the experimental results clearly support this point. To avoid bias that may arise from adding input only between user/project tokens, we introduced a comparison with fixed additive attention. We found that simply adding fixed connection information to attention calculation for nodes in the graph is not effective. It is truly effective to include information that reflects the actual relationships between nodes.
@@ 4.2.5. Study of Parameters
This experiment aims to answer: Can we ensure consistency between our pre-training and fine-tuning tasks? We conducted experiments on the AM-Toys dataset to analyze the performance alignment between the pre-training task and the fine-tuning task. We used the results of the first 10 pre-training epochs and the corresponding loss function. Then, we fine-tuned the pre-trained model to obtain evaluation metrics. We compared the 3 metrics, Recall@20, Recall@40, and NDCG@100 with the loss function. Figure  6 shows the trend of changes in the 3 metrics is consistent with the trend of changes in the loss functions. This indicates that our pre-training task and fine-tuning task are well-aligned, and our prompt construction method can provide rich information for subsequent recommendation tasks.
@@ 5. Related Work
The application of LLMs in recommendation systems mainly contains two types: deep representation of data using LLMs, and direct application of generative LLMs to construct recommendation logic.
@@
In deep representation, discriminative language models like BERT are widely used for fine-tuning and pre-training, integrating specific domain data features to enhance the performance of recommendation systems. For instance, U-BERT (Qiu et al., 2021) leverages content-rich domain data to learn user representations, compensating for the scarcity of behavioral data. Similarly, UserBERT (Wu et al., 2021) includes two self-supervised tasks for pretraining on unlabeled behavior data. Additionally, BECR (Yang et al., 2022) combines deep contextual token interactions with traditional lexical word matching features. Notably, the ”pretrain-finetune” mechanism plays a crucial role in sequence or session-based recommendation systems, like BERT4Rec (Sun et al., 2019) and RESETBERT4Rec (Zhao, 2022). UniSRec (Hou et al., 2022) develops a BERT fine-tuning framework that links item description texts. In content-based recommendations, especially in the news domain, models like NRMS (Wu et al., 2021), Tiny-NewsRec (Yu et al., 2021), and PREC (Liu et al., 2022) enhance news recommendations by leveraging LLMs, addressing domain transfer issues or reducing transfer costs. Research by Penha and Hauff (Penha and Hauff, 2020) shows that BERT, even without fine-tuning, effectively prioritizes relevant items in ranking processes, illustrating the potential of large language models in natural language understanding.
@@
In recent studies, generative LLMs have shown huge potential in recommendation systems through prompting and tuning methods. Notable works and advancements include: Liu et al. (Liu et al., 2023a) conducted a comprehensive assessment of ChatGPT’s performance in five key recommendation tasks. Sanner et al. (Sanner et al., 2023) designed three different prompt templates to evaluate the enhancement effect of prompts, finding that zero-shot and few-shot strategies are particularly effective in preference-based recommendations using language. Sileo et al. (Sileo et al., 2022) and Hou et al. (Hou et al., 2023) focused on designing effective prompt methods for specific recommendation tasks. Gao and team (Gao et al., 2023) developed ChatREC around ChatGPT, an interactive recommendation framework that understands user needs through multiple rounds of dialogue. Petrov and Macdonald (Petrov and Macdonald, 2023) introduced GPTRec, a generative sequence recommendation model based on GPT-2. Kang and colleagues (Kang et al., 2023) explored formatting user historical interactions as prompts and assessed the performance of LLMs of different scales. Dai et al. (Dai et al., 2023) designed templates for various recommendation tasks using demonstration example templates. Bao et al. (Bao et al., 2023) developed TALLRec, which demonstrates the potential of LLMs in recommendation domains through two-stage fine-tuning training. Ji et al. (Ji et al., 2023) presented GenRec, a method that leverages the generative capabilities of LLMs to directly generate the target of recommendations. In specific scenarios like online recruitment, generative recommendation models such as GIRL (Zheng et al., 2023) and reclm (Friedman et al., 2023) demonstrated enhanced explainability and appropriateness in recommendations. Li et al. (2023e) (Li et al., 2023) described user behaviors and designed prompts in news with PBNR. Wang et al. (Penha and Hauff, 2020) proposed UniCRS, a design based on knowledge-enhanced rapid learning.

### v2

@@ 4.2.2. Ablation Studies
  • LLM-NoGHIP does not include graph or historical information to embed prompts for pre-training but only uses simple users’ review information to evaluate the impact of complex prompts on model performance.
We pre-trained and fine-tuned each baseline model separately, and then compared it with our complete model. These pre-training and fine-tuning experimental settings are consistent and conducted on the same dataset to ensure the comparability of results. Table 4 shows each component’s specific contribution to the model’s overall performance. For example, the performance of LLM-NoPretrain is significantly lower than that of the complete model. This implies that using recommendation-related graph data and natural language data for pre-training plays a crucial role in improving model performance. Similarly, the results of LLM-NoFineTune demonstrate the importance of fine-tuning. Subsequently, by comparing the performance of LLM-NoGKIA, and LLM-NoGHIP with that of the complete model, we find that the addition of graph connection information in attention calculation and complex prompts containing second-order relationships is crucial for improving the performance of recommender systems.
Table 4. Ablation Study Results.
Dataset Metric LLM-NoPretrain LLM-NoFineTune LLM-NoGKIA LLM-NoGHIP Ours
AM-Beauty Recall@20 0.0464 0.0441 0.1225 0.1267 0.1590
Recall@40 0.0709 0.0691 0.1665 0.1799 0.2177
NDCG@100 0.0339 0.0323 0.0790 0.0827 0.1029
AM-Toys Recall@20 0.0477 0.0580 0.0896 0.0858 0.1349
Recall@40 0.0689 0.1003 0.1272 0.1179 0.1873
NDCG@100 0.0330 0.0481 0.0612 0.0594 0.0876
AM-Sports Recall@20 0.0449 0.0394 0.0555 0.0558 0.0764
Recall@40 0.0719 0.0613 0.0846 0.0830 0.1240
NDCG@100 0.0322 0.0278 0.0391 0.0379 0.0535
AM-Luxury Recall@20 0.1872 0.1885 0.2474 0.2679 0.3066
Recall@40 0.2233 0.2254 0.2880 0.3028 0.3441
NDCG@100 0.1223 0.1235 0.1834 0.2065 0.2331
AM-Scientific Recall@20 0.0708 0.0668 0.1383 0.1206 0.1480
Recall@40 0.1037 0.0960 0.1822 0.1575 0.1908
NDCG@100 0.0568 0.0465 0.0940 0.0810 0.1072
AM-Instruments Recall@20 0.0766 0.0727 0.1387 0.1426 0.1698
Recall@40 0.1004 0.0948 0.1741 0.1779 0.2265
NDCG@100 0.0500 0.0478 0.1042 0.1044 0.1312
AM-Food Recall@20 0.0224 0.0204 0.1275 0.1264 0.1438
Recall@40 0.0299 0.0274 0.1559 0.1487 0.1673
NDCG@100 0.0153 0.0141 0.0898 0.0963 0.1119
4.2.3. Study on Different Pre-training Prompt Structures
This experiment aims to answer: What is the contribution of pre-training text data that integrates second-order relationships and item background information to recommendation models? To answer this question, we chose the AM-Toys dataset and designed the following experimental models:
@@ 4.2.3. Study on Different Pre-training Prompt Structures
  • Without Prompt: The model does not contain any second-order relationship information or item background information.
@@
Figure  4 shows that the model that uses prompt sentences of complete information (with the second-order relationship) performs best over all the performance indicators. The performances of the ”Without second-order relationship” model are lower than that of the complete model. As can be seen, second-order relationship information is an essential component of graph connectivity. Similarly, the ”Without Item” model performs poorly, highlighting the importance of natural language background information in enhancing recommender systems.
@@
4.2.4. Study on Different Historical Interaction Length
This experiment aims to answer: How does the historical interaction length influence the recommendation performance? Here, we employ different numbers of historical interactions in the fine-tuning prompts. Because in real-world scenarios, there would be numerous historical interactions. Selecting appropriate numbers will find a balance between cache memory and performance.
Figure 5. Results of Different Historical Length.
The results show that the length of historical events does affect the performance. However, the difference between different parameters is not significant. Therefore, when employing the system, we have more flexibility to choose the appropriate historical length.
@@ 4.2.5. Study on Different Attention Injection Ways
Fig. 6 shows that the model using our graph attentive LLM method exhibits the best performance. Our method not only considers the direct connections between nodes but also the spatial relationships (i.e., the shortest connected path) between nodes in the graph.
Figure 7. Results of Two Parts of $R$.
We compared our method with regular attention mechanisms, and the experimental results clearly support this point. To avoid bias that may arise from adding input only between user/project tokens, we introduced a comparison with fixed additive attention. We found that simply adding fixed connection information to attention calculation for nodes in the graph is not effective. It is truly effective to include information that reflects the actual relationships between nodes. Meanwhile, our experiments on the AM-Toys dataset, with results shown in Fig. 7 show that both direct $R^{\text{conn}}$ and indirect connective information $R^{\text{path}}$ contribute to the performance.
@@ 4.2.6. Study of Parameters
This experiment aims to answer: Can we ensure consistency between our pre-training and fine-tuning tasks? We conducted experiments on the AM-Toys dataset to analyze the performance alignment between the pre-training task and the fine-tuning task. We used the results of the first 10 pre-training epochs and the corresponding loss function. Then, we fine-tuned the pre-trained model to obtain evaluation metrics. We compared the 3 metrics, Recall@20, Recall@40, and NDCG@100 with the loss function. Fig. 8 shows the trend of changes in the 3 metrics is consistent with the trend of changes in loss functions. This indicates that our pre-training task and fine-tuning task are well-aligned, and our prompt construction method can provide rich information for subsequent recommendation tasks.
@@
4.2.7. Study of Different LLM backbones
To explore the generalizability and robustness of our method across different large language model (LLM) architectures, we further conduct experiments using BART as the backbone, in addition to GPT-2. Unlike GPT-2, which is a decoder-only transformer, BART is a sequence-to-sequence model that integrates both transformer encoder and decoder components. This architectural difference allows BART to potentially better capture bidirectional contextual information, which may influence recommendation outcomes.
(a) Standard Attention
(b) Graph Information Integrated
Figure 9. Comparison between Different LLM Backbones.
As shown in Fig. 9, we compare the performance of GPT and BART under two settings: (a) with standard attention, and (b) with graph information integration. The models are evaluated using NDCG@100, Recall@40, and Recall@20.
From the results, we observe that:
  • Under standard attention, the BART-based model performs marginally better than GPT in all three metrics.
  • However, with graph information integrated, the performance gap diminishes, and GPT slightly outperforms BART in Recall@20.
These findings indicate that while BART’s encoder-decoder structure brings benefits in general scenarios, GPT-based models remain competitive and may perform better when enhanced with graph-structured information. This highlights the adaptability of our framework across different LLMs and the importance of backbone selection based on task characteristics.
4.2.8. Study of Cold-start Situation
To evaluate the performance of our model in cold-start scenarios, we conduct additional experiments by simulating users with limited interaction history. Specifically, we randomly reduce each user’s training item list by 50%, which will limit the available data for LLM to learn user preferences. This setup mimics the cold-start condition where users have interacted with only a few items.
Table 5. Performance under Normal and Cold-start Settings.
Dataset Metric BERT4Rec-Normal BERT4Rec-Cold RecMind-Normal RecMind-Cold Ours-Normal Ours-Cold
AM-Beauty Recall@20 0.1126 0.0921 0.1347 0.1206 0.1590 0.1384
Recall@40 0.1677 0.1439 0.1874 0.1634 0.2177 0.1944
NDCG@100 0.0781 0.0631 0.0846 0.0730 0.1029 0.0812
AM-Toys Recall@20 0.0853 0.0711 0.1126 0.0999 0.1349 0.1274
Recall@40 0.1375 0.1181 0.1564 0.1392 0.1873 0.1763
NDCG@100 0.0532 0.0435 0.0584 0.0506 0.0876 0.0654
AM-Sports Recall@20 0.0521 0.0451 0.0683 0.0605 0.0764 0.0658
Recall@40 0.0701 0.0584 0.1147 0.1020 0.1240 0.1102
NDCG@100 0.0305 0.0258 0.0511 0.0452 0.0535 0.0502
AM-Luxury Recall@20 0.2076 0.1735 0.2879 0.2499 0.3066 0.3044
Recall@40 0.2404 0.1968 0.3351 0.2860 0.3441 0.3389
NDCG@100 0.1617 0.1393 0.2049 0.1832 0.2331 0.2005
AM-Scientific Recall@20 0.0871 0.0734 0.1274 0.1112 0.1480 0.1445
Recall@40 0.1160 0.0989 0.1651 0.1427 0.1908 0.1876
NDCG@100 0.0606 0.0517 0.0873 0.0774 0.1072 0.1023
AM-Instruments Recall@20 0.1183 0.1009 0.1539 0.1359 0.1698 0.1544
Recall@40 0.1531 0.1268 0.2093 0.1877 0.2265 0.1943
NDCG@100 0.0922 0.0761 0.1008 0.0851 0.1312 0.1163
AM-Food Recall@20 0.1036 0.0863 0.1194 0.1073 0.1438 0.1316
Recall@40 0.1284 0.1130 0.1399 0.1262 0.1673 0.1546
NDCG@100 0.0835 0.0713 0.0783 0.0680 0.1119 0.0996
Table 5 shows the results, where we compare the performance of our method with baseline models under both normal and cold-start settings. Our model demonstrates relatively stable performance degradation, suggesting its robustness in data-sparse scenarios. Compared to BERT4Rec and RecMind, our method consistently achieves the best performance across all datasets in both settings. Notably, although all models experience performance drops when user interaction data is limited, the decline in our method is significantly smaller, especially on complex datasets such as AM-Luxury and AM-Scientific. For example, on AM-Luxury, our method only drops 0.0022 in Recall@20, whereas BERT4Rec drops 0.0341 and RecMind drops 0.0380. This indicates that our model can better capture user intent with limited supervision. We attribute this advantage to the dual-view representation and the semantic richness preserved in our LLM-based design, which allows our model to generalize better under cold-start conditions. These results further confirm the effectiveness of our approach in improving recommendation robustness in real-world scenarios.
4.2.9. Scalability Analysis
To better understand the practical feasibility of our proposed model, we conduct additional experiments to evaluate its training and inference efficiency under different data scales. We select three Amazon datasets of varying sizes—AM-Luxury (small), AM-Toys (medium), and AM-Food (large)—to simulate realistic scenarios with increasing user-item interaction volumes.
We measure and report the following metrics:
  • Training time per epoch (in seconds)
  • Validation inference time (in seconds)
  • Peak GPU memory usage (in MB)
The results are summarized in Table 6. The training and inference times grow approximately linearly with the dataset size, while the peak memory consumption remains relatively stable across different datasets. These findings demonstrate that our model is computationally efficient and scalable to larger recommendation tasks.
We also observe that the inference time is longer than the training time per epoch. This is because, during inference, the model must compute scores for all candidate items for each user and perform a complete ranking. In contrast, training only requires calculating the loss based on sampled interactions. The full ranking process increases the computational workload and inference latency.
Table 6. Training and Inference Efficiency on Datasets of Different Sizes.
Dataset Training Time (s/epoch) Inference Time (s) Peak Memory (MB)
AM-Luxury 4.46 5.02 5002.08
AM-Toys 38.14 51.13 20018.38
AM-Food 233.85 327.18 26096.70
Overall, these experimental results verify that the proposed graph-attentive LLM recommender maintains good computational efficiency and can be effectively deployed in large-scale recommendation systems.
4.2.10. Case Study: Graph-Aware Recommendation Reasoning
We present a case study based on a single user from the AM-Toys dataset to demonstrate how our model leverages structural connectivity between users and items during recommendation.
Figure 10. A case study showing the model’s recommendation results and their connections to the user’s historical interactions.
As illustrated in Figure 10, the model receives a prompt indicating that user_5546 has previously interacted with four items: item_1105, item_537, item_1904, and item_485. Based on this interaction history, the model is prompted to generate a personalized recommendation list.
The predicted list includes items such as item_3875, item_2464, and item_2401. Notably, item_3875 appears in the user’s actual future interactions, confirming the correctness of the prediction. Even for items not directly interacted with, we observe meaningful structural or semantic connections. For instance, item_1464 shares the same category as item_1108, and item_2401 belongs to the same brand as item_1108. This case highlights how the model captures both direct and indirect connections in the interaction graph to generate meaningful and interpretable recommendations.
@@ 5. Related Work
Recommender systems (RS) play a crucial role in various applications, helping users navigate vast amounts of information efficiently. Traditional RS methods have been extensively studied and widely applied, laying a strong foundation for the field. Recently, the emergence of LLMs has introduced new possibilities, broadly categorized into two approaches: (1) leveraging LLMs for deep representation learning to enhance user and item embeddings and (2) directly applying generative LLMs to construct recommendation logic, enabling more adaptive and context-aware recommendations.
5.1. Traditional Recommender System
Traditional recommendation algorithms began with similarity-based methods, particularly Collaborative Filtering (CF), which exploits user-item interactions (Koren et al., 2009; Salakhutdinov and Mnih, 2008; Rendle et al., 2009). Early CF models, such as Matrix Factorization (MF) (Koren et al., 2009), effectively captured latent preferences but struggled with data sparsity and linearity constraints. The rise of neural networks introduced models like Neural Collaborative Filtering (He et al., 2017), which replaced inner product-based interactions with multi-layer perceptrons for greater expressiveness. More recently, Graph Neural Networks (GNNs) have become prominent in recommendation  (Wang et al., 2019; He et al., 2020; Bai et al., 2024a). Neural Graph Collaborative Filtering (Wang et al., 2019) and LightGCN (He et al., 2020) propagate user-item interaction signals through graph structures, effectively capturing high-order relationships. Beyond model architecture, research has expanded into diversified open problems, such as multi-modal recommendation, leveraging textual, visual, and knowledge graph data (He and McAuley, 2016; He et al., 2024a); denoising-based recommendation, addressing noisy interactions (He et al., 2024b); debias recommendation, tackling issues like popularity bias (Cai et al., 2024b, a); cold-start recommendation, recommending to new user without historical interactions (Bai et al., 2023, 2024b). Another direction is feature engineering, which enhances recommendation performance by refining input representations (Lian et al., 2018; Lin et al., 2022). Feature selection techniques identify the most informative features to reduce noise and improve generalization (Ying et al., 2024b; Wang et al., 2024; Gong et al., 2025e; Ying et al., 2025b, 2024c), while feature transformation, such as autoencoders and reinforcement learning-based approaches, enable the construction of more expressive feature spaces (Ying et al., 2024d, 2023; Gong et al., 2025c; Wang et al., 2025b; Ying et al., 2024a; Gong et al., 2025d, b; Hu et al., 2024; bai2025privacy,wang2025llm). These methods contribute to the robustness and adaptability of recommendation systems across various domains. This evolution from CF to neural and graph-based models, alongside advances in feature engineering and diverse research topics, reflects the ongoing advancement in recommender systems.
5.2. Discriminative LLMs for Recommender System
@@
In deep representation, discriminative language models like BERT are widely used for fine-tuning and pre-training, integrating specific domain data features to enhance the performance of recommender systems. For instance, U-BERT (Qiu et al., 2021) leverages content-rich domain data to learn user representations, compensating for the scarcity of behavioral data. Similarly, UserBERT (Wu et al., 2021) includes two self-supervised tasks for pretraining on unlabeled behavior data. Additionally, BECR (Yang et al., 2022) combines deep contextual token interactions with traditional lexical word matching features. Notably, the ”pretrain-finetune” mechanism plays a crucial role in sequence or session-based recommender systems, like BERT4Rec (Sun et al., 2019) and RESETBERT4Rec (Zhao, 2022). UniSRec (Hou et al., 2022) develops a BERT fine-tuning framework that links item description texts. In content-based recommendations, especially in the news domain, models like NRMS (Wu et al., 2021), Tiny-NewsRec (Yu et al., 2021), and PREC (Liu et al., 2022) enhance news recommendations by leveraging LLMs, addressing domain transfer issues, or reducing transfer costs. Research by Penha and Hauff (Penha and Hauff, 2020) shows that BERT, even without fine-tuning, effectively prioritizes relevant items in ranking processes, illustrating the potential of large language models in natural language understanding. DWSRec (Zhang et al., 2024) relies only on pre-trained text embeddings without the need for ID embeddings. By applying dual-view whitening, it enhances the isotropy of embeddings while preserving semantic information. LLMRec (Wei et al., 2024) enhances recommender systems by leveraging LLMs for graph-based augmentation and denoised data robustification, effectively mitigating data sparsity. CoLLM (Zhang et al., 2025) integrates collaborative filtering information as a separate modality into large language models (LLMs) via an external mapping module, aligning collaborative embeddings with text inputs.
5.3. Generative LLMs for Recommender System
@@
Recent advances in generative models have combined neural generation with symbolic reasoning, enabling more interpretable and structured decision-making (Ying et al., 2025a; Gong et al., 2025a). Building on this, generative LLMs have shown strong potential in recommender systems through prompting, fine-tuning, and routing (Wang et al., 2025a). Notable works and advancements include: Liu et al. (Liu et al., 2023a) conducted a comprehensive assessment of ChatGPT’s performance in five key recommendation tasks. Sanner et al. (Sanner et al., 2023) designed three different prompt templates to evaluate the enhancement effect of prompts, finding that zero-shot and few-shot strategies are particularly effective in preference-based recommendations using language. Sileo et al. (Sileo et al., 2022) and Hou et al. (Hou et al., 2023) focused on designing effective prompt methods for specific recommendation tasks. Gao and team (Gao et al., 2023) developed ChatREC around ChatGPT, an interactive recommendation framework that understands user needs through multiple rounds of dialogue. Petrov and Macdonald (Petrov and Macdonald, 2023) introduced GPTRec, a generative sequence recommendation model based on GPT-2. Kang and colleagues (Kang et al., 2023) explored formatting user historical interactions as prompts and assessed the performance of LLMs of different scales. PageLLM (Wang et al., 2025c) introduces a multi-grained reward model to fine-tune the LLM using reinforcement learning from human feedback. Dai et al. (Dai et al., 2023) designed templates for various recommendation tasks using demonstration example templates. Bao et al. (Bao et al., 2023) developed TALLRec, which demonstrates the potential of LLMs in recommendation domains through two-stage fine-tuning training. Ji et al. (Ji et al., 2023) presented GenRec, a method that leverages the generative capabilities of LLMs to directly generate the target of recommendations. In specific scenarios like online recruitment, generative recommendation models such as GIRL (Zheng et al., 2023) and reclm (Friedman et al., 2023) demonstrated enhanced explainability and appropriateness in recommendations. Li et al.  (Li et al., 2023) described user behaviors and designed prompts in news with PBNR. Wang et al. (Wang et al., 2022) proposed UniCRS, a design based on knowledge-enhanced rapid learning. HSTU (Zhai et al., 2024) is a generative recommendation model that treats recommendation as a sequence generation task, achieving high accuracy and efficiency on large-scale recommendation data. RecMind (Wang et al., 2023) is an LLM-based recommendation agent that makes zero-shot personalized recommendations by using external tools and a self-inspiring planning method to better use past information.